The principal claims that the students in his school are above average intelligence. A random sample of 35 IQ scores have a mean score of 110.5. Is there sufficient evidence to support the principal’s claim? The mean population IQ is 100 with a standard deviation of 25. Use =1%.


1
Expert's answer
2022-06-07T15:59:15-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le100"

"H_1:\mu>100"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z:z>2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{110.5-100}{25/\sqrt{35}}=2.4848"

Since it is observed that "z=2.4848>2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z>2.4848)=0.006481," and since "p=0.006481<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 100, at the "\alpha = 0.01" significance level.